### v1

# Divergence-Based Similarity Function for Multi-View Contrastive Learning

Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of an instance. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across various tasks, including kNN classification and linear evaluation, while also offering greater efficiency compared to other multi-view methods. Furthermore, we establish a theoretical connection between DSF and cosine similarity, and show that, unlike cosine similarity, DSF operates effectively without requiring a temperature hyperparameter.
1 Introduction
Figure 1: Illustration of feature‐distribution modeling on the unit hypersphere: given multiple augmented views of the image, their extracted feature vectors (dots) lie on the hyperspherical manifold and are used to construct smooth probability distributions (colored regions) over the sphere.
Learning effective representations without costly labels is a central challenge in modern machine learning. Self-supervised learning (SSL) addresses this by deriving supervisory signals from the data itself, and contrastive learning has emerged as one of its most successful paradigms. In contrastive frameworks such as SimCLR[Chen et al., 2020a], MoCo[He et al., 2020], and DCL[Yeh et al., 2022], representations of two augmented views of the same instance are ”pulled together” while being ”pushed apart” from all other samples. A key ingredient across these methods is cosine similarity on the unit hypersphere, which normalizes features to lie on a sphere and then measures their dot-product similarity.
Figure 2: Illustration of multi-view contrastive learning strategies. A data instance is augmented $M$ times, resulting in $x^{1},\cdots,x^{M}$. Base encoder outputs $M$ representations $z^{1},\cdots,z^{M}$ from $M$ views. From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views before applying contrast; and (iii) DSF, our proposed model that aggregates the view’s feature into distribution and fuses these distributions into a single contrastive loss
However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches.:
  • Loss-level approaches: methods that formulate the contrastive objective by averaging pairwise losses computed between all available combinations of views, typically using arithmetic or geometric averaging schemes, as in [Shidani et al., 2024].
  • Feature-level approaches: methods that first aggregate multi-view representations—most notably through feature averaging—before applying a contrastive objective. For example, FastSiam [Pototzky et al., 2022] computes the average of view features, while other techniques like M3G[Piran et al., 2024] adopt more complex strategies such as multi-marginal transport.
Although these multi-view extensions generalize two-view objectives, they still rely solely on pairwise relations and overlook the joint structural relationships among the full set of views.
To overcome these limitations, we introduce a novel similarity function that directly models the joint structure among all $M$ augmented views of an instance. Rather than reducing multi-view interactions to pairwise comparisons or expensive transport computations, our method first estimates a probability distribution over the views on the unit hypersphere and then computes the similarity between two instances via a negative divergence measure between their respective distributions. In particular, we employ the von Mises–Fisher distribution, which is parameterized by a mean direction $\mu$ and a concentration parameter $\kappa$. The mean direction $\mu$ represents the central tendency of the distribution, while the concentration parameter $\kappa$ captures the joint structure among the multi-view features. Moreover, under the special case of two views, our divergence based similarity exactly recovers the cosine similarity used in the InfoNCE loss, thereby bridging the gap between conventional pairwise and multi-view contrastive paradigms.
Our main contributions are:
  • Divergence-based multi-view Contrastive Learning. we propose a Divergence-based Similarity Function(DSF) that captures joint structure across all views without decomposing into pairwise terms.
  • Relation with original InfoNCE loss. We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering a unified view of pairwise and multi-view contrastive methods.
  • Hyperparameter-free temperature control We mathematically establish the role of the temperature term in contrastive loss and show that our model learns optimally without manual temperature adjustment.
  • Experimental Results. We validate the effectiveness of our method through extensive experiments on kNN classification, linear evaluation, and resource analysis in terms of GPU memory and training time, consistently outperforming the state-of-the-art contrastive learning approaches.
The remainder of this paper is structured as follows. Section 2 provides an overview of contrastive learning and its multi-view extensions. Section 3 reviews related work on contrastive frameworks with both two-view ($M=2$) and multi-view ($M\geq 3$) settings. In Section 4, we present our proposed divergence-based similarity formulation. Section 5 describes the experimental setup and results. Finally, Section 6 concludes the paper and discusses potential directions for future research.
2 Preliminaries
(a) kNN Top-1 Accuracy with respect to training time.
(b) Linear evaluation accuracy with respect to training time.
Figure 3: More Time Cost Comparison of performance over training time for kNN and linear evaluation. Both figures report accuracy for different training epochs and wall-clock time under a fixed GPU memory budget, with the number of views set to 8 ($M=2m=8$). The right-hand side of each figure shows the elapsed time of MoCo baselines, while the x-axis indicates the time and epochs for the multi-view methods. Note that the lowest dashed line (800 epochs, 5.5h) corresponds to the same GPU memory and training time as the first point(250 epochs, 5.5h) of each multi-view method. Compared to Loss_avg and Feature_avg at 800 epochs, our proposed DSF achieves more than 2$\times$ faster convergence in the kNN setting and over 3$\times$ faster convergence in the linear evaluation setting.
2.1 Contrastive Learning
Contrastive Learning (CL) has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs. Here, positives refer to semantically similar pairs whereas negatives to dissimilar ones. The core idea is to map positive samples closer in the embedding space while pushing negative samples apart. Given a positive pair and a set of negative samples, this loss encourages the model to assign high similarity to the positive pair relative to the negatives. The loss can be formulated as follows:
$L_{i}=-\log\frac{e^{(sim(z_{i},z^{+}_{i})/\tau)}}{e^{(sim(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{i\neq j}^{K}e^{(sim(z_{i},z^{-}_{j})/\tau)}}$ (1)
where $z^{+}_{i}$ is the embeddings of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of $K-1$ negative pairs. $sim(\cdot,\cdot$) is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
2.2 von Mises-Fisher Distribution
The von Mises-Fisher (vMF) distribution is a probability distribution dealing with a data that lies on a unit hypersphere. Defined over a $(p-1)$-dimensional unit hypersphere $\mathbb{S}_{p-1}$, vMF distribution describes the probability density of a unit vector with a specified mean direction. The distribution is characterized by a mean direction, indicating the central orientation of the vector, and a concentration parameter, which describes how tightly the distribution clusters around this mean direction.
The probability density function of the vMF distribution for a unit vector $x\in S_{p-1}$ is given by:
$f_{p}(x;,\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x)$ (2)
where $\mu$ is the mean direction, $\kappa$ is the concentration parameter, and $C_{p}(\kappa)$ is a normalization constant.
A higher concentration parameter $\kappa$ results in samples that are tightly clustered around the mean direction $\mu$, while $\kappa=0$ corresponds to a uniform distribution over the hypersphere.
3 Related Works
3.1 Original Contrastive Learning
Contrastive Learning appears in many self-supervised learning methods. SimCLR[Chen et al., 2020a] leverages the InfoNCE objective[Oord et al., 2018] to pull together representations of two augmented views while pushing apart all other samples. MoCo[He et al., 2020] maintains a momentum encoder, which is a moving average of the parameters to supply a large, consistent set of negatives without labels. [Yeh et al., 2022] decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable, high performance training with smaller batches and without a momentum encoder. MoCo v3 [Chen et al., 2021] explores contrastive learning using a transformer-based model, specifically the Vision Transformer (ViT) [Dosovitskiy et al., 2020].
3.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view contrastive learning methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. [Shidani et al., 2024] generate $M$ views per sample and aggregate their pairwise losses either arithmetically or geometrically. The second category is a method that deal with view representations at the feature level before computing the final objective. FastSiam[Pototzky et al., 2022] averages the other $M-1$ view features to define each view’s target , and M3G[Piran et al., 2024] adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, our method integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
4 Divergence-based similarity function
In this section, we explain the details of proposed divergence-based similarity function(DSF) and relations with cosine similarity function. In Sec4.1, we formalize notations throughtout this paper. In Sec4.2, we describe how to estimate vMF distribution parameters $\mu,\kappa$. In Sec4.3,we propose the adjustment to kappa value for computational stabilization. In Sec4.4, we describe the calculation of DSF. In Sec4.5, we show that KL divergence-based similarity function between vMF distribution becomes equivalent to consine similarity function in InfoNCE Loss under certain conditions. Furthermore, we show why our divergence similairty function does not need a temperature hyperparameter and why cosine similarity function need temperature hyperparamter.
4.1 Notation
We denote the batch size, number of views, and feature dimensionality by $B$, $M$, and $D$, respectively. Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a von Mises–Fisher (vMF) distribution by the calligraphic symbol $\mathcal{D}$.
Let $x$ be an image instance, and $x^{l}$ denote its $l$-th augmented view. Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$.
contrastive Learning(CL) makes $\mathbb{R}^{B\times 2\times D}$ features with two augmentation with each instance. Multi-view contrastive learning makes $\mathbb{R}^{B\times 2m\times D}$ features with total $M=2m$ views with each instance.
In CL, there are $2B$ features $z_{i}$, $i=1,\cdots,2B$. $z_{i}$ and $z_{B+i}$ are positively correlated. In Multi-view CL, there are $B\cdot 2m$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$. $\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
Dataset Base Model Method Epoch kNN↑ Linear Evaluation↑
ImageNet-100 MoCo v3 - 300 74.08 81.52
MoCo v3 loss_avg 100 76.62 84.24
MoCo v3 fea_avg 100 76.94 84.06
MoCo v3 DSF 100 79.00 85.66
CIFAR10 MoCo - 800 88.44 88.10
MoCo loss_avg 250 88.28 88.58
MoCo fea_avg 250 87.47 88.02
MoCo DSF 250 90.04 91.21
Table 1: Comparison of kNN and Linear Evaluation Results Under Equivalent Resource Settings. All experiments are conducted under the same GPU memory and time constraints. The number of views is fixed at $M=2m=8$.
4.2 Estimate vMF Distribution
Before introducing the DSF, we first describe the process of constructing von Mises–Fisher (vMF) distributions from multiple augmented views of a single image instance.
Given an image instance $x$, we generate $M$ augmented views via data augmentation. These views are divided into two groups of $m$ samples each. All augmented images are passed through the encoder to obtain $M$ feature vectors on the unit hypersphere. Each group of features is then used to estimate the parameters of a vMF distribution. As a result, each instance is represented by two vMF distributions, each constructed from one group of augmented features.
Let $z^{1},\dots,z^{m}$ denote the feature vectors corresponding to the $m$ augmented views of an instance $x$. The mean direction $\mu$ and concentration parameter $\kappa$ of the vMF distribution are estimated as follows:
$\mu=\frac{\bar{z}}{\bar{R}},\quad\kappa=A_{p}^{-1}(\bar{R}),$ (3)
where
$\bar{z}=\frac{1}{m}\sum\limits_{l=1}^{m}z^{l},\quad\bar{R}=\|\bar{z}\|.$ (4)
Since computing $\kappa=A_{p}^{-1}(\bar{R})$ is intractable in practice, we instead adopt an approximate form of $\kappa$ proposed by [Banerjee et al., 2005]:
$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}$ (5)
Consequently, the vMF distribution constructed from the features $z^{1},\dots,z^{m}$ is denoted as:
$\mathcal{D}(z^{1},\dots,z^{m})=\mathcal{D}(\mu,\kappa).$ (6)
For subscript indexing $i$ :
$\mathcal{D}(z_{i}^{1},\dots,z_{i}^{m})=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
4.3 Stabilization of the Concentration Parameter $\kappa$
As a final step before introducing the DSF, we describe a strategy to stabilize the concentration parameter $\kappa$, which is essential for maintaining numerical stability during training.
As training progresses, the value of $\kappa$ can grow excessively, often leading to numerical overflow in computations. This issue arises from two primary factors:
  • High feature dimensionality. For a fixed mean resultant length $\bar{R}$, a larger feature dimension $p$ results in a sharper increase in $\kappa$. This becomes problematic, particularly when using higher-dimensional feature spaces (e.g., $D=128$ in our setup), as commonly done in larger models.
  • Feature concentration. As the encoder learns more discriminative representations, the extracted features become increasingly concentrated, causing $\bar{R}$ to approach 1, which in turn further amplifies $\kappa$.
Although this behavior is theoretically valid, it leads to numerical instability when computing divergence measures in practice. To mitigate this issue, we apply two simple yet effective techniques:
  1. 1.
    Dimension-based normalization: We normalize $\kappa$ by the feature dimension $p$, as $\kappa$ grows approximately linearly with $p$ when the mean resultant length $\bar{R}$ is fixed, as shown in Eq. 5. This prevents dimension-driven growth.
  2. 2.
    Scaling of mean resultant length: To avoid $\bar{R}$ approaching 1, we apply a scaling factor $\lambda_{\bar{R}}\in(0,1)$. In our experiments, we use $\lambda_{\bar{R}}=0.95$, modifying $\bar{R}$ as follows:
    $\bar{R}\longrightarrow\lambda_{\bar{R}}\cdot\bar{R}$ (8)
These modifications are empirically shown to stabilize training and avoid divergence explosion in high-dimensional settings. Further details and illustrative figures can be found in the Appendix C Figure 6.
4.4 Divergence-based similarity function
4.4.1 Similarity function
Similarity function $\text{sim}(\cdot,\cdot)$ is given by $\text{sim}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. For example, cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$. Accordingly, DSF handles multi-view features, and its similarity function is defined as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$.
4.4.2 KL Divergence of vMF Distribution
Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distribution. The KL divergence value of $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms.
$\begin{split}&D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})\\ &=\log\frac{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{i},\kappa_{i})}{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{j},\kappa_{j})}\\ &=\left(\frac{p}{2}-1\right)\log\frac{\kappa_{i}}{\kappa_{j}}+\log\frac{I_{\frac{p}{2}-1}(\kappa_{j})}{I_{\frac{p}{2}-1}(\kappa_{i})}\\ &\qquad\qquad\qquad+A_{p}(\kappa_{i})\cdot\left(\kappa_{i}-\kappa_{j}\cdot\mu_{i}^{T}\mu_{j}\right)\\ \\ \end{split}$ (9)
The KL divergence range in from 0 to infinity. KL divergence is not a distance function of two distribution. As the KL divergence value close to zero, it means the distributions are similar i.e. $\mu,\kappa$ become close. So, our proposed divergence-based similarity function is negative KL divergence value as following.
$\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})=-D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})$ (10)
$L_{i,\text{InfoNCE}}^{\text{div}}=-\log\left(\frac{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}}{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})}}\right)$ (11)
Here, $\mathcal{D}_{i}^{+}$ denotes the positive distribution paired with $\mathcal{D}_{i}$. The negative distributions are denoted by $\mathcal{D}_{j}$, where $K$ represents the number of negative distributions. Note that the value of $K$ may vary depending on the training framework.
4.5 Analysis : Relation with Cosine Similarity Function
Next, we establish the connection between the conventional cosine similarity function and our proposed DSF. We further explain why cosine similarity requires a temperature scaling term for effective optimization, whereas our DSF inherently eliminates the need for such adjustment.
Theorem 1 (Relation between Divergence-based and Cosine Similarities).
Let $Z_{i}=\{z^{1}_{i},\cdots,z^{m}_{i}\}$ and $Z_{j}=\{z^{1}_{j},\cdots,z^{m}_{j}\}$ are $m$ multi-view features. $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ are the estimated von Mises-Fisher distribution from $Z_{i}$ and $Z_{j}$. Let $\text{sim}_{\text{cos}}(\cdot,\cdot)$ is cosine similarity and $\text{sim}_{\text{div}}(\cdot,\cdot)$ is the negative Kullback-Leibler divergence between two vMF distribution. If m=1, $\kappa_{1}$=$\kappa_{2}$ and $A_{p}(\kappa_{1})\cdot\kappa_{1}=1/\tau$ then
$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\text{sim}_{\text{cos}}(z_{i},z_{j})/\tau+constant$ (12)
Furthermore, InfoNCE Loss with $sim_{\text{div}}$ becomes equivalent to that using $sim_{\text{cos}}$ one:
$L^{\text{div}}_{\text{InfoNCE}}=L^{\text{cos}}_{\text{InfoNCE}}$ (13)
Proof.
Proof is at Appendix A. ∎
Theorem 1 says that when we have two views and kappa is set under certain conditions. Then our proposed divergence-based simiarity function becomes cosine simiarity function with some constant value. Furthermore, this leads to fact that InfoNCE loss becomes equivalent.
Now, we explain the dependency of cosine similarity on temperature hyperparameter and independence of our DSF with respect to temperature hyperparameter.
Proposition 1.
Let $K$ be the number of negative pairs, and assume all negative similarities have the same value. Let $s^{+}$ and $s^{-}$ denote the positive and negative similarities, respectively. Then, the InfoNCE loss becomes:
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(s^{+}-s^{-})}{\exp(s^{+}-s^{-})+K}\\ (s^{+}-s^{-})&\in\left\{\begin{matrix}[-\frac{2}{\tau},\frac{2}{\tau}]&\text{if }\text{sim}=\text{sim}_{\text{cos}}\\ (-\infty,\infty)&\text{if }\text{sim}=\text{sim}_{\text{div}}\end{matrix}\right.\end{split}$ (14)
In the optimal case for cosine similarity, we have $s^{+}=+1$ and $s^{-}=-1$, resulting in:
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(2/\tau)}{\exp(2/\tau)+K}\\ &=6.3196,\quad\tau=1.0,\,K=4096\\ &=4.3310,\quad\tau=0.5,\,K=4096\\ &=0.1706,\quad\tau=0.2,\,K=4096\\ &=0.0000,\quad\tau=0.1,\,K=4096\\ \end{split}$ (15)
It has been theoretically shown (see Proposition1) that cosine similarity cannot represent a appropriate loss value without the use of a temperature term $\tau$. The core issue arises from the limited range of the similarity margin $(s^{+}-s^{-})$, which restricts the contrast between positive and negative pairs. To compensate for this limitation, the temperature parameter $\tau$ is introduced to amplify this margin, allowing for better gradient signals during training.
The number of negative samples $K$ varies depending on the contrastive learning framework. For instance, in SimCLR and DCL, $K=2B-2=254$ when the batch size $B=128$, while in MoCo, $K$ is determined by the size of the memory queue, ranging from 4,096 to 65,536 depending on the dataset.
As summarized in Table 2, the InfoNCE loss yields inappropriate values without the temperature parameter $\tau$ in all $K$ values. It says cosine similarity requires a carefully tuned temperature parameter across all frameworks to function effectively. In contrast, our proposed DSF expresses an appropriate loss without the need for such a temperature hyperparameter. See Figure 5 for $(s^{+}-s^{-})$ margin in our DSF experiment.
$L_{\text{InfoNCE}}$
$\tau$ K=256 K=4096 K=65536
1.0 3.573 6.319 9.090
0.5 1.738 4.331 7.091
0.2 0.011 0.170 1.380
0.1 0.000 0.000 0.0001
Table 2: InfoNCE loss values under optimal similarity ($s^{+}=1$, $s^{-}=-1$) for varying temperature $\tau$ and the number of negative pairs $K$.
Algorithm 1 Three Multi-View Methods
Input: Instance $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
Parameter: Number of views $M=2m$.
Output: Divergence-based similarity between positive pairs.
1:  Apply $M=2m$ augmentations to $x_{i}$, yielding $x_{i}^{1},\dots,x_{i}^{m}$ and $x_{B+i}^{1},\dots,x_{B+i}^{m}$
2:  Extract features:
$z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\quad z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\quad$
for $\text{for }l=1,\dots,m$
3:  if multi-view method = loss_avg:
$L_{\text{loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
4:  if multi-view method = fea_avg:
$z_{i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$
$L_{\text{fea\_avg}}=L_{\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
5:  if multi-view method = DSF:  Estimate vMF distribution parameters as in Eq. (3)–(4):
$(\mu_{i},\kappa_{i})\leftarrow\{z_{i}^{1},\dots,z_{i}^{m}\}$
$(\mu_{B+i},\kappa_{B+i})\leftarrow\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}$
6:   Compute divergence-based similarity:
$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$
$L_{\text{DSF}}=L_{\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
5 Experiments
5.1 Experimental Setup
Every experiment is conducted with one or four NVIDIA GeForce RTX 3090.
Dataset
This study utilizes two benchmark datasets: CIFAR-10 and ImageNet-100. CIFAR-10 [Krizhevsky et al., 2009] consists of 60,000 color images of size $32\times 32$ pixels, divided into 50,000 training and 10,000 test images across 10 object classes. ImageNet-100 is a curated subset of the ImageNet dataset, comprising 100 randomly selected classes with approximately 1,300 images per class. It serves as a mid-scale benchmark for evaluating representation learning performance.
These datasets are used to validate the proposed method on both small-scale and mid-scale visual recognition tasks.
Model Arichitecture
We primarily follow the MoCo v2 [Chen et al., 2020b] framework for contrastive learning. Throughout this paper, any reference to ”MoCo” refers to MoCo v2 unless otherwise specified. We adopt ResNet-18 as the encoder for experiments on CIFAR-10 The final fully connected layer projects features into a 128-dimensional embedding space.
For large-scale evaluation, we extend our experiments to ImageNet-100 using MoCo v3 [Chen et al., 2021] with a Vision Transformer (ViT-S/16) backbone. This enables assessment of the scalability of the proposed method across both convolutional and transformer-based architectures.
Data Augmentation
For data augmentation, we apply a combination of RandAugment [Cubuk et al., 2020], ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in contrastive learning setups.
5.2 Evaluations
To evaluate the model’s performance, we conducted following tasks: k-Nearest Neighbors (kNN) and Linear Evaluation. Linear Evaluation assesses the representational quality of the encoder by attaching a classifier, while kNN measures similarity-based performance directly within the encoder’s embedding space.
kNN Evaluation.
We evaluated the model with a kNN approach to measure the encoder’s capacity to retain similarity in the embedding space. Evaluation was conducted on CIFAR-10, ImageNet-100 using the output feature of the trained query encoder. The K value was set to 200, enabling an assessment of the encoder’s representations in a non-parametric, unsupervised manner. In our experiments, the proposed model achieved a kNN accuracy of 79.00% on the ImageNet-100 dataset, outperforming the baseline MoCo v3 model, which attained 74.08%. On the CIFAR-10 dataset, our model achieved 90.04% accuracy, compared to 88.44% with the MoCo baseline.
Linear Evaluation.
To further validate the effectiveness of our method, we evaluate the quality of the learned representations through linear evaluation. Specifically, after performing self-supervised pretraining on ImageNet-100 and CIFAR-10 using our proposed DSF framework, we freeze the encoder and train a linear classifier (a single-layer MLP) on top of the extracted features.
On the ImageNet-100 dataset, our method achieved a linear evaluation accuracy of 85.66%, outperforming the baseline MoCo v3 model, which attained 81.52%. Similarly, on the CIFAR-10 dataset, our method achieved an accuracy of 91.21%, compared to 88.10% obtained by the MoCo baseline. These results demonstrate that DSF leads to stronger and more transferable representations across different datasets.
(a) kNN Top-1 Accuracy with respect to number of view.
(b) Linear Evaluation Top-1 Accuracy with respect to number of view.
Figure 4: Comparison of kNN and linear evaluation accuracies according to the number of views in three multi-view learning methods.
5.3 Same Resource Comparison
Our method samples $M$ multiple augmented views from each image, which could potentially increase computational overhead. Therefore, it is essential to ensure that the proposed approach remains efficient under equivalent computational constraints.
Computational Efficiency.
Our method samples $M$ augmented views from each image instance, which could potentially increase computational overhead. To address this, we ensure fair comparisons by maintaining consistent computational budgets across all experiments.
To ensure the practicality of our method, we carefully control and match computational budgets across all experiments. Specifically, we fix the product of batch size and number of views—i.e., $B\times M$—so that memory usage remains constant regardless of the number of views. For example, in our ImageNet-100 experiments (see Table 1), the baseline MoCo is trained with a configuration of $512\times 2$, whereas our method and other multi-view baselines use $128\times 8$, maintaining identical GPU memory usage. This configuration strategy is also consistently applied to CIFAR-10 experiments.
Moreover, to equalize the overall training time, we proportionally reduce the number of training epochs for multi-view methods. As a result, the total runtime difference between our approach and baseline models is strictly limited to within 30 minutes—less than 4% and 10% of the entire training duration in the ImageNet-100 and CIFAR-10 setting. These measures ensure that the observed performance improvements are attributable to our proposed DSF method itself, rather than to increased computational resources.
5.4 More Resource Comparison
5.4.1 More Time Cost: Faster Convergence with Divergence Similarity
Self-supervised learning methods are known to require longer training times for convergence due to the nature of instance discrimination objectives and the inherent challenge of learning meaningful representations from unlabeled data. This issue becomes even more pronounced when dealing with complex datasets or limited batch sizes.
To investigate convergence behavior, we extended the number of training epochs and compared model accuracy as a function of elapsed training time, as shown in Figure 3. While the baseline MoCo gradually improves over a long training schedule (55 hours at 8000 epochs), our proposed DSF reaches comparable accuracy within only 11 hours at 400 epochs, using merely 20% of the total training time. Specifically, compared to the Loss_avg and Feature_avg baselines trained for 800 epochs, our DSF reaches similar accuracy within just 400 epochs, achieving a 2× speedup in convergence. Moreover, we observe that the performance gap between our DSF and the two other multi-view baselines (Loss_avg, Feature_avg) remains consistently large, regardless of the total training time. This indicates that our method not only converges faster but also achieves a higher final accuracy that is not easily matched even with extended training.
This result highlights that our divergence-based formulation facilitates more efficient optimization in self-supervised learning. Notably, all three multi-view methods considered in this comparison have identical per-epoch training time, ensuring a fair evaluation.
5.4.2 More GPU Memory : Number of Views
We sample multiple augmented views from each image; thus, it is important to investigate how the number of sampled views affects the quality of the learned representations. Prior studies [Tian et al., 2020, Shidani et al., 2024, Pototzky et al., 2022] have shown that increasing the number of augmentations generally leads to better representation quality.
In our framework, the number of views directly impacts the accuracy of the estimated von Mises–Fisher (vMF) distribution. As the number of views increases, the estimated parameters $\mu$ and $\kappa$ become more reliable due to reduced estimation variance.
Figure 4 presents the performance trends of three multi-view learning methods on both kNN and linear evaluation tasks, as the number of views $M$ increases. Across all view counts, DSF consistently outperforms the other two approaches. These results suggest that our divergence-based formulation provides stable and reliable improvements in multi-view learning settings.
5.5 Temperature Free
In numerous studies, the temperature parameter $\tau$ has been shown to be critical for achieving optimal performance in contrastive learning [Zhang et al., 2021, Zhang et al., 2022, Kukleva et al., 2023]. In our method, however, $\tau$ is not required, as discussed in Proposition 1, and it can even degrade performance when introduced.
To validate this, we conducted experiments over a wide range of temperature values, from 0.1 to 5.0. Interestingly, setting $\tau=1$ (i.e., using the raw similarity without scaling) consistently resulted in the best performance in Table 3. This suggests that our DSF is inherently numerically stable and does not require additional temperature tuning.
Temperature $\tau$ 0.8 0.9 1 1.1
kNN 91.2 91.3 91.8 91.4
Linear Evaluation 92.0 92.3 92.5 92.0
Table 3: Temperature Experiment. MoCo + DSF results in CIFAR10.
5.6 Training Details
To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch. We use AdamW as our optimizer. The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder. We choose batch size of 512, learning rate of 0.03. In our proposed method, a larger learning rate of 0.06 is used. GPU used for overall training is NVIDIA GeForce RTX 3090.
6 Future Work
Figure 5: Training dynamics of the similarity margin $(s^{+}-s^{-})$. Values are averaged over mini-batches throughout training. This plot shows the similarity margin of MoCo v3 with DSF ($128\times 8$) configuration on ImageNet-100.
We use von Mises-Fisher distribution and Kullback-Leibler divergence for our proposed divergence-based similarity function. Although they are popular choice, there are plenty of available distributions and divergences. For example, in divergence, there are other divergences such as Rényi divergence, $\chi^{\alpha}$divergence, and Jensen–Shannon divergence—collectively known as the family of $f$-divergences could be explored. Investigating how these alternative divergences influence the representation quality and computational efficiency of contrastive methods remains a promising direction for future work.
Acknowledgements
This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127) and MSIT/IITP(2021-0-00077)
Appendix A Appendix: Proof of Theorem 1
We prove that the divergence-based similarity function becomes equivalent to the cosine similarity function in InfoNCE loss under specific conditions.
Assume $m=1$, $\kappa_{1}=\kappa_{2}=\kappa$, and $A_{p}(\kappa)\cdot\kappa=1/\tau$. Then, the divergence-based similarity simplifies as:
$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})+C,$ (16)
where $C$ is a constant independent of $(z_{i},z_{j})$ pair.
Substituting this into the InfoNCE loss:
$\begin{split}&L_{i,\text{InfoNCE}}^{\text{div}}\\ &=-\log\left(\frac{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}}{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})}}\right)\\ &=-\log\left(\frac{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})+C}}{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})+C}+\sum\limits_{j\neq i}^{K}e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})+C}}\right)\\ &=-\log\left(\frac{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})}}{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})}}\right)\\ &=L_{i,\text{InfoNCE}}^{\text{cos}}.\end{split}$ (17)
Appendix B Feature Averaging Also Encodes Only Pairwise Relationships
Loss_avg is clearly a pairwise-based method. We further demonstrate that Feature_avg, despite aggregating multiple views, still considers only pairwise relationships among features, as can be shown through a straightforward calculation.
$z_{i}^{\text{avg}}$ $=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$
$\text{sim}_{\text{cos}}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$ $=(z_{i}^{\text{avg}})^{T}z_{B+i}^{\text{avg}}$
$=\left(\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l}\right)^{T}\left(\frac{1}{m}\sum_{l^{\prime}=1}^{m}z_{B+i}^{l^{\prime}}\right)$
$=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}(z_{i}^{l})^{T}z_{B+i}^{l^{\prime}}$ (18)
As shown above, the final similarity score is an unweighted average over all pairwise cosine similarities between individual view features $z_{i}^{l}$ and $z_{B+i}^{l^{\prime}}$. Thus, Feature_avg does not model joint structure beyond pairwise relationships.
Appendix C Stabilization of the Concentration Parameter $\kappa$
To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Equation 5. However, even this approximation can lead to excessively large values of $\kappa$ as training progresses. To mitigate this, we propose two stabilization techniques.
Figure 6 illustrates the original approximated $\kappa$ values computed from Equation 5, alongside the stabilized values obtained after applying our proposed techniques.
Figure 6: Consistency by Kappa Normalization. $\kappa$ value is calculated as approximated formulation as (5). $x$-axis is $\bar{R}$ and $y$-axis is calculated $\kappa$. $p$ is feature dimension size. In our backbone ResNet18, feature dimension is 128. (upper) Directly calculated $\kappa$. (lower) $\kappa$ is divided by feature dimension $p$ and calculated with $\bar{R}$ multiplied by $\lambda_{\bar{R}}$. Without these two techniques, $\kappa$ value is going excessively large.

### v2

@@
 Abstract
-Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of an instance.
+Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of data.
 While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views.
 In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions.
-Extensive experiments demonstrate that DSF consistently improves performance across various tasks, including kNN classification and linear evaluation, while also offering greater efficiency compared to other multi-view methods.
-Furthermore, we establish a theoretical connection between DSF and cosine similarity, and show that, unlike cosine similarity, DSF operates effectively without requiring a temperature hyperparameter.
+Extensive experiments demonstrate that DSF consistently improves performance across diverse tasks—including kNN classification, linear evaluation, transfer learning, and distribution shift—while also achieving greater efficiency than other multi-view methods.
+Furthermore, we establish a connection between DSF and cosine similarity, and demonstrate that, unlike cosine similarity, DSF operates effectively without the need for tuning a temperature hyperparameter.
 1 Introduction
-Figure 1: Illustration of feature‐distribution modeling on the unit hypersphere: given multiple augmented views of the image, their extracted feature vectors (dots) lie on the hyperspherical manifold and are used to construct smooth probability distributions (colored regions) over the sphere.
-Learning effective representations without costly labels is a central challenge in modern machine learning.
-Self-supervised learning (SSL) addresses this by deriving supervisory signals from the data itself, and contrastive learning has emerged as one of its most successful paradigms.
-In contrastive frameworks such as SimCLR[Chen et al., 2020a], MoCo[He et al., 2020], and DCL[Yeh et al., 2022], representations of two augmented views of the same instance are ”pulled together” while being ”pushed apart” from all other samples.
-A key ingredient across these methods is cosine similarity on the unit hypersphere, which normalizes features to lie on a sphere and then measures their dot-product similarity.
+Self-supervised learning (SSL) aims to learn meaningful representations without labels, often using contrastive learning (CL) frameworks such as SimCLR (Chen et al.
+2020a), MoCo (He et al.
+2020), and DCL (Yeh et al.
+2022).
+These methods encourage similarity between augmented views of the same data while separating different data, typically via cosine similarity on unit-normalized features.
+However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches:
+  • Loss-level approaches: Averaging pairwise losses computed between all available combinations of views (Shidani et al. 2024).
+  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky, Sultan, and Schmidt-Thieme 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
+Figure 1: Illustration of divergence-based similarity function: given multiple augmented views of the image, their extracted feature vectors (colored dots) lie on the hypersphere and are used to construct probability distributions (colored regions) over the unit hypersphere.
 Figure 2: Illustration of multi-view contrastive learning strategies.
-A data instance is augmented $M$ times, resulting in $x^{1},\cdots,x^{M}$.
+Given an input $x$, we generate $M$ augmented views, resulting in $x^{1},\cdots,x^{M}$.
 Base encoder outputs $M$ representations $z^{1},\cdots,z^{M}$ from $M$ views.
-From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views before applying contrast; and (iii) DSF, our proposed model that aggregates the view’s feature into distribution and fuses these distributions into a single contrastive loss
-However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches.:
-  • Loss-level approaches: methods that formulate the contrastive objective by averaging pairwise losses computed between all available combinations of views, typically using arithmetic or geometric averaging schemes, as in [Shidani et al., 2024].
-  • Feature-level approaches: methods that first aggregate multi-view representations—most notably through feature averaging—before applying a contrastive objective.
-For example, FastSiam [Pototzky et al., 2022] computes the average of view features, while other techniques like M3G[Piran et al., 2024] adopt more complex strategies such as multi-marginal transport.
+From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views; and (iii) DSF, our proposed method, which aggregates features from $M$ multiple views into distributions and computes a single contrastive loss between them.
 Although these multi-view extensions generalize two-view objectives, they still rely solely on pairwise relations and overlook the joint structural relationships among the full set of views.
-To overcome these limitations, we introduce a novel similarity function that directly models the joint structure among all $M$ augmented views of an instance.
-Rather than reducing multi-view interactions to pairwise comparisons or expensive transport computations, our method first estimates a probability distribution over the views on the unit hypersphere and then computes the similarity between two instances via a negative divergence measure between their respective distributions.
-In particular, we employ the von Mises–Fisher distribution, which is parameterized by a mean direction $\mu$ and a concentration parameter $\kappa$.
-The mean direction $\mu$ represents the central tendency of the distribution, while the concentration parameter $\kappa$ captures the joint structure among the multi-view features.
-Moreover, under the special case of two views, our divergence based similarity exactly recovers the cosine similarity used in the InfoNCE loss, thereby bridging the gap between conventional pairwise and multi-view contrastive paradigms.
+To address this, we propose a novel similarity function that models the joint structure among all augmented views of data.
+We estimate a probability distribution on the unit hypersphere and compute similarity via negative divergence between distributions, see Figure 1.
+We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher, Lewis, and Embleton 1993; Kitagawa and Rowley 2022; Scott, Gallagher, and Mozer 2021; Du et al.
+2024).
+vMF distribution is parameterized by mean direction $\mu$ and concentration $\kappa$, which respectively represent the central tendency and joint structure of features.
+Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord, Li, and Vinyals 2018), linking pairwise and multi-view contrastive frameworks.
 Our main contributions are:
   • Divergence-based multi-view Contrastive Learning. We propose a Divergence-based Similarity Function (DSF) that captures joint structure across all views without decomposing into pairwise terms.
-  • Relation with original InfoNCE loss.
-We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering a unified view of pairwise and multi-view contrastive methods.
-  • Hyperparameter-free temperature control We mathematically establish the role of the temperature term in contrastive loss and show that our model learns optimally without manual temperature adjustment.
-  • Experimental Results.
-We validate the effectiveness of our method through extensive experiments on kNN classification, linear evaluation, and resource analysis in terms of GPU memory and training time, consistently outperforming the state-of-the-art contrastive learning approaches.
+  • Establishing the relation to the original InfoNCE loss.
+We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering the unified view of pairwise and multi-view contrastive methods.
+  • Temperature-free. We show the dependency of the temperature term in cosine similarity and show that, for DSF, temperature tuning is unnecessary.
+  • Superior Experimental Results.
+We validate the effectiveness of DSF through extensive experiments on kNN classification, linear evaluation, transfer learning, and distribution shift, as well as resource analysis in terms of GPU memory and training time, consistently outperforming both baseline and existing multi-view contrastive learning methods.
 The remainder of this paper is structured as follows.
-Section 2 provides an overview of contrastive learning and its multi-view extensions.
 Section 2 reviews related work on contrastive frameworks with both two-view ($M=2$) and multi-view ($M\geq 3$) settings.
-In Section 4, we present our proposed divergence-based similarity formulation.
+Section 3 provides an overview of CL and its multi-view extensions.
+Section 4 presents our proposed formulation.
 Section 5 describes the experimental setup and results.
 Finally, Section 6 concludes the paper and discusses potential directions for future research.
-2 Preliminaries
+2 Related Works
 (a) kNN Top-1 Accuracy with respect to training time.
 (b) Linear evaluation accuracy with respect to training time.
-Figure 3: More Time Cost Comparison of performance over training time for kNN and linear evaluation.
-Both figures report accuracy for different training epochs and wall-clock time under a fixed GPU memory budget, with the number of views set to 8 ($M=2m=8$).
-The right-hand side of each figure shows the elapsed time of MoCo baselines, while the x-axis indicates the time and epochs for the multi-view methods.
-Note that the lowest dashed line (800 epochs, 5.5h) corresponds to the same GPU memory and training time as the first point(250 epochs, 5.5h) of each multi-view method.
-Compared to Loss_avg and Feature_avg at 800 epochs, our proposed DSF achieves more than 2$\times$ faster convergence in the kNN setting and over 3$\times$ faster convergence in the linear evaluation setting.
+[…]
+2.1 Original Contrastive Learning
+CL appears in many self-supervised learning methods.
+SimCLR (Chen et al.
+2020a) leverages the InfoNCE loss (Oord, Li, and Vinyals 2018) to pull together representations of two augmented views while pushing apart all other samples.
+MoCo (He et al.
+2020) introduces a momentum encoder and a dynamic dictionary (negative queue) to maintain a large and consistent set of negative samples.
+DCL (Yeh et al.
+2022) decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable and high performance training with smaller batches and without a momentum encoder.
+MoCo v3 (Chen, Xie, and He 2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al.
+2020).
+2.2 Contrastive Learning with $M\geq 3$ views
+Existing $M\geq 3$ multi‐view CL methods have broadly fallen into two categories.
+The first category shapes the loss by averaging pairwise terms across views.
+(Shidani et al.
+2024) generates $M$ views per a sample and aggregates their pairwise losses either arithmetically or geometrically.
+The second category is a method that deals with view representations at the feature level before computing the final objective.
+FastSiam (Pototzky, Sultan, and Schmidt-Thieme 2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al.
+2024) adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost.
+In contrast to these approaches, DSF integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
+3 Preliminaries
 3.1 Contrastive Learning
-Contrastive Learning (CL) has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs.
-Here, positives refer to semantically similar pairs whereas negatives to dissimilar ones.
-The core idea is to map positive samples closer in the embedding space while pushing negative samples apart.
-Given a positive pair and a set of negative samples, this loss encourages the model to assign high similarity to the positive pair relative to the negatives.
+CL has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs.
+Positive pair refers to semantically similar pair whereas negative pair to dissimilar ones.
+The core idea is to map positive samples closer while pushing negative samples apart in the embedding space.
+Given a positive pair and a set of negative samples, InfoNCE loss encourages the model to assign high similarity to the positive pair relative to the negatives.
 The loss can be formulated as follows:
-$L_{i}=-\log\frac{e^{(sim(z_{i},z^{+}_{i})/\tau)}}{e^{(sim(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{i\neq j}^{K}e^{(sim(z_{i},z^{-}_{j})/\tau)}}$ (1)
-where $z^{+}_{i}$ is the embeddings of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of $K-1$ negative pairs.
-$sim(\cdot,\cdot$) is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
+$L^{\mathrm{cos}}_{i,\text{InfoNCE}}=-\log\frac{e^{(\mathrm{sim}(z_{i},z^{+}_{i})/\tau)}}{e^{(\mathrm{sim}(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{j\neq i}^{K}e^{(\mathrm{sim}(z_{i},z^{-}_{j})/\tau)}},$ (1)
+where $z^{+}_{i}$ is the embedding of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of K1K-1 negative pairs.
+$\mathrm{sim}(\cdot,\cdot)$ is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
 3.2 von Mises-Fisher Distribution
-The von Mises-Fisher (vMF) distribution is a probability distribution dealing with a data that lies on a unit hypersphere.
-Defined over a $(p-1)$-dimensional unit hypersphere $\mathbb{S}_{p-1}$, vMF distribution describes the probability density of a unit vector with a specified mean direction.
-The distribution is characterized by a mean direction, indicating the central orientation of the vector, and a concentration parameter, which describes how tightly the distribution clusters around this mean direction.
-The probability density function of the vMF distribution for a unit vector $x\in S_{p-1}$ is given by:
-$f_{p}(x;,\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x)$ (2)
-where $\mu$ is the mean direction, $\kappa$ is the concentration parameter, and $C_{p}(\kappa)$ is a normalization constant.
-A higher concentration parameter $\kappa$ results in samples that are tightly clustered around the mean direction $\mu$, while $\kappa=0$ corresponds to a uniform distribution over the hypersphere.
-3 Related Works
-3.1 Original Contrastive Learning
-Contrastive Learning appears in many self-supervised learning methods.
-SimCLR[Chen et al., 2020a] leverages the InfoNCE objective[Oord et al., 2018] to pull together representations of two augmented views while pushing apart all other samples.
-MoCo[He et al., 2020] maintains a momentum encoder, which is a moving average of the parameters to supply a large, consistent set of negatives without labels.
-[Yeh et al., 2022] decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable, high performance training with smaller batches and without a momentum encoder.
-MoCo v3 [Chen et al., 2021] explores contrastive learning using a transformer-based model, specifically the Vision Transformer (ViT) [Dosovitskiy et al., 2020].
-3.2 Contrastive Learning with $M\geq 3$ views
-Existing $M\geq 3$ multi‐view contrastive learning methods have broadly fallen into two categories.
-The first category shapes the loss by averaging pairwise terms across views.
-[Shidani et al., 2024] generate $M$ views per sample and aggregate their pairwise losses either arithmetically or geometrically.
-The second category is a method that deal with view representations at the feature level before computing the final objective.
-FastSiam[Pototzky et al., 2022] averages the other $M-1$ view features to define each view’s target , and M3G[Piran et al., 2024] adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost.
-In contrast to these approaches, our method integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
+vMF distribution is defined on the (p1)(p\!-\!1)-dimensional unit hypersphere $\mathbb{S}_{p-1}$ and models the probability density of unit vectors with a specified mean direction $\mu$ and concentration parameter $\kappa$.
+Its probability density function for a unit vector $x\in\mathbb{S}_{p-1}$ is:
+$f_{p}(x;\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x),$ (2)
+where $C_{p}(\kappa)$ is the normalization constant.
+Higher $\kappa$ values yield distributions more concentrated around $\mu$, while $\kappa=0$ corresponds to a uniform distribution on the unit hypersphere.
+3.3 Notation
+We denote the batch size, number of views, and feature dimension by $B$, $M$, and $p$, respectively.
+Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
+Let $x$ be an image data, and $x^{l}$ is denoted by its $l$-th augmented view.
+Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$.
+In CL, there are $2\cdot B$ features $z_{i}$, $i=1,\cdots,2B$.
+$z_{i}$ and $z_{B+i}$ are positively correlated.
+In multi-view CL, there are $2m\cdot B$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$.
+$\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
+The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a vMF distribution by the calligraphic symbol $\mathcal{D}$.
 4 Divergence-based similarity function
-In this section, we explain the details of proposed divergence-based similarity function(DSF) and relations with cosine similarity function.
-In Sec4.1, we formalize notations throughtout this paper.
-In Sec4.2, we describe how to estimate vMF distribution parameters $\mu,\kappa$.
-In Sec4.3,we propose the adjustment to kappa value for computational stabilization.
-In Sec4.4, we describe the calculation of DSF.
-In Sec4.5, we show that KL divergence-based similarity function between vMF distribution becomes equivalent to consine similarity function in InfoNCE Loss under certain conditions.
-Furthermore, we show why our divergence similairty function does not need a temperature hyperparameter and why cosine similarity function need temperature hyperparamter.
-4.1 Notation
-We denote the batch size, number of views, and feature dimensionality by $B$, $M$, and $D$, respectively.
-Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
-The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a von Mises–Fisher (vMF) distribution by the calligraphic symbol $\mathcal{D}$.
-Let $x$ be an image instance, and $x^{l}$ denote its $l$-th augmented view.
-Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$.
-contrastive Learning(CL) makes $\mathbb{R}^{B\times 2\times D}$ features with two augmentation with each instance.
-Multi-view contrastive learning makes $\mathbb{R}^{B\times 2m\times D}$ features with total $M=2m$ views with each instance.
-In CL, there are $2B$ features $z_{i}$, $i=1,\cdots,2B$.
-$z_{i}$ and $z_{B+i}$ are positively correlated.
-In Multi-view CL, there are $B\cdot 2m$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$.
-$\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
-Dataset Base Model Method Epoch kNN↑ Linear Evaluation↑
-ImageNet-100 MoCo v3 - 300 74.08 81.52
-MoCo v3 loss_avg 100 76.62 84.24
-MoCo v3 fea_avg 100 76.94 84.06
-MoCo v3 DSF 100 79.00 85.66
-CIFAR10 MoCo - 800 88.44 88.10
-MoCo loss_avg 250 88.28 88.58
-MoCo fea_avg 250 87.47 88.02
-MoCo DSF 250 90.04 91.21
-Table 1: Comparison of kNN and Linear Evaluation Results Under Equivalent Resource Settings.
-All experiments are conducted under the same GPU memory and time constraints.
-The number of views is fixed at $M=2m=8$.
-4.2 Estimate vMF Distribution
-Before introducing the DSF, we first describe the process of constructing von Mises–Fisher (vMF) distributions from multiple augmented views of a single image instance.
-Given an image instance $x$, we generate $M$ augmented views via data augmentation.
-These views are divided into two groups of $m$ samples each.
-All augmented images are passed through the encoder to obtain $M$ feature vectors on the unit hypersphere.
-Each group of features is then used to estimate the parameters of a vMF distribution.
-As a result, each instance is represented by two vMF distributions, each constructed from one group of augmented features.
-Let $z^{1},\dots,z^{m}$ denote the feature vectors corresponding to the $m$ augmented views of an instance $x$.
-The mean direction $\mu$ and concentration parameter $\kappa$ of the vMF distribution are estimated as follows:
+In this section, we describe DSF in detail and demonstrate its relationship to cosine similarity, as well as its inherent temperature-free property.
+4.1 Estimation of vMF Distribution
+First, we describe how to construct vMF distributions from multiple augmented views of an image.
+Given data $x$, we sample $M$ augmented views and divide them into two groups of $m$ samples each.
+These are passed through the encoder to obtain unit-normalized feature vectors.
+Each group is then used to estimate a vMF distribution, yielding two vMFs per a data.
+Let $z^{1},\dots,z^{m}$ be the feature vectors from one group. The mean direction $\mu$ and concentration $\kappa$ are estimated as:
 $\mu=\frac{\bar{z}}{\bar{R}},\quad\kappa=A_{p}^{-1}(\bar{R}),$ (3)
@@
 $\bar{z}=\frac{1}{m}\sum\limits_{l=1}^{m}z^{l},\quad\bar{R}=\|\bar{z}\|.$ (4)
-Since computing $\kappa=A_{p}^{-1}(\bar{R})$ is intractable in practice, we instead adopt an approximate form of $\kappa$ proposed by [Banerjee et al., 2005]:
-$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}$ (5)
-Consequently, the vMF distribution constructed from the features $z^{1},\dots,z^{m}$ is denoted as:
-$\mathcal{D}(z^{1},\dots,z^{m})=\mathcal{D}(\mu,\kappa).$ (6)
-For subscript indexing $i$ :
-$\mathcal{D}(z_{i}^{1},\dots,z_{i}^{m})=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
+As $A_{p}^{-1}$ is intractable, we use the approximation from (Banerjee et al. 2005):
+$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}.$ (5)
+Let $g_{\text{est}}$ be the estimation function mapping features to vMF parameters:
+$g_{\text{est}}(z^{1},\dots,z^{m})=(\mu,\kappa),$
+and define the resulting vMF distribution as:
+$\mathcal{D}\left(g_{\text{est}}(z^{1},\dots,z^{m})\right)=\mathcal{D}(\mu,\kappa).$ (6)
+For data $i$, we denote:
+$\mathcal{D}\left(g_{\text{est}}(z_{i}^{1},\dots,z_{i}^{m})\right)=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
 4.2 Stabilization of the Concentration Parameter $\kappa$
-As a final step before introducing the DSF, we describe a strategy to stabilize the concentration parameter $\kappa$, which is essential for maintaining numerical stability during training.
-As training progresses, the value of $\kappa$ can grow excessively, often leading to numerical overflow in computations. This issue arises from two primary factors:
-  • High feature dimensionality.
-For a fixed mean resultant length $\bar{R}$, a larger feature dimension $p$ results in a sharper increase in $\kappa$.
-This becomes problematic, particularly when using higher-dimensional feature spaces (e.g., $D=128$ in our setup), as commonly done in larger models.
-  • Feature concentration.
-As the encoder learns more discriminative representations, the extracted features become increasingly concentrated, causing $\bar{R}$ to approach 1, which in turn further amplifies $\kappa$.
-Although this behavior is theoretically valid, it leads to numerical instability when computing divergence measures in practice. To mitigate this issue, we apply two simple yet effective techniques:
+In this section, we describe a stabilization strategy for the concentration parameter $\kappa$, which is critical for preventing numerical issues during training.
+As training progresses, $\kappa$ may grow excessively due to:
+  • High feature dimensionality: Larger feature dimensions $p$ result in sharper increase in $\kappa$, especially with high-dimensional embeddings (e.g., $D=128$).
+The problem is further exacerbated when larger models are used.
+  • Feature concentration: As representations improve, features become more aligned and $\bar{R}$ approaches 1, further sharp increasing $\kappa$.
+Though theoretically valid, such growth causes numerical instability in divergence computation. To mitigate this, we adopt two strategies:
   1. 1.
-    Dimension-based normalization: We normalize $\kappa$ by the feature dimension $p$, as $\kappa$ grows approximately linearly with $p$ when the mean resultant length $\bar{R}$ is fixed, as shown in Eq.
-5.
-This prevents dimension-driven growth.
+    Dimension-based normalization: Normalize $\kappa$ by the feature dimension $p$, since $\kappa$ grows roughly linearly with respect to $p$ when $\bar{R}$ is fixed (see Eq. 5).
   2. 2.
-    Scaling of mean resultant length: To avoid $\bar{R}$ approaching 1, we apply a scaling factor $\lambda_{\bar{R}}\in(0,1)$.
-In our experiments, we use $\lambda_{\bar{R}}=0.95$, modifying $\bar{R}$ as follows:
+    Scaling mean resultant length: Prevent $\bar{R}$ from approaching 1 by applying a scaling factor $\lambda_{\bar{R}}=0.95$:
     $\bar{R}\longrightarrow\lambda_{\bar{R}}\cdot\bar{R}$ (8)
-These modifications are empirically shown to stabilize training and avoid divergence explosion in high-dimensional settings.
-Further details and illustrative figures can be found in the Appendix C Figure 6.
+These techniques are applied after estimating the vMF parameters, and they help stabilize training and prevent divergence-related issues in high-dimensional settings.
+Further details are provided in Appendix C.3, Figure 6.
+Dataset Base Model Method Epoch kNN (↑) Linear Evaluation (↑) mCE* (↓)
+ImageNet-100 MoCo v3 - 300 74.08 83.42 36.67
+MoCo v3 Loss_avg 100 76.62 84.06 38.27
+MoCo v3 Feature_avg 100 76.94 83.82 39.42
+MoCo v3 DSF 100 79.00 85.80 33.15
+CIFAR-10 MoCo - 800 88.44 88.10 19.03
+MoCo Loss_avg 250 88.28 88.58 19.90
+MoCo Featrue_avg 250 87.47 88.02 19.80
+MoCo DSF 250 90.04 91.21 16.20
+Table 1: Comparison of kNN, Linear Evaluation and mCE Results Under Equivalent Resource Settings.
+All experiments are conducted under the same GPU memory and time constraints.
+The number of views is fixed at $M=2m=8$.
+The mean Corruption Error(mCE) was evaluated on CIFAR-10-C and ImageNet-100-C, with all models trained solely on clean(uncorrupted) data.
 4.3 Divergence-based similarity function
-4.4.1 Similarity function
+Similarity function
 Similarity function $\text{sim}(\cdot,\cdot)$ is given by $\text{sim}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$.
-For example, cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$.
-Accordingly, DSF handles multi-view features, and its similarity function is defined as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$.
-4.4.2 KL Divergence of vMF Distribution
-Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distribution.
-The KL divergence value of $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms.
-$\begin{split}&D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})\\ &=\log\frac{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{i},\kappa_{i})}{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{j},\kappa_{j})}\\ &=\left(\frac{p}{2}-1\right)\log\frac{\kappa_{i}}{\kappa_{j}}+\log\frac{I_{\frac{p}{2}-1}(\kappa_{j})}{I_{\frac{p}{2}-1}(\kappa_{i})}\\ &\qquad\qquad\qquad+A_{p}(\kappa_{i})\cdot\left(\kappa_{i}-\kappa_{j}\cdot\mu_{i}^{T}\mu_{j}\right)\\ \\ \end{split}$ (9)
-The KL divergence range in from 0 to infinity.
-KL divergence is not a distance function of two distribution.
-As the KL divergence value close to zero, it means the distributions are similar i.e.
-$\mu,\kappa$ become close.
-So, our proposed divergence-based similarity function is negative KL divergence value as following.
+For example, the cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$.
+Accordingly, we define the DSF similarity function to handle multi-view features as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$.
+More precisely,
+[…]
+KL Divergence of vMF Distribution
+Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distributions.
+The KL divergence value between $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms:
+[…]
+Figure 7 in the Appendix illustrates the behavior of each term as well as the overall KL divergence values.
+Detailed explanation and evaluation of $A_{p}(\cdot)$ and $I_{p}(\cdot)$ are provided in Appendix C.
+The KL divergence ranges from 0 to infinity.
+As the KL divergence approaches zero, the two distributions become similar, indicating that $\mu_{i}\approx\mu_{j}$ and $\kappa_{i}\approx\kappa_{j}$.
+Conversely, a large KL divergence implies that the two distributions are highly dissimilar.
+Therefore, we define the DSF as the negative KL divergence.
+We use following expression for notational convenience throughout the remainder of the paper:
 $\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})=-D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})$ (11)
+Now, we define the loss function with DSF as:
 $L_{i,\text{InfoNCE}}^{\text{div}}=-\log\left(\frac{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}}{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})}}\right)$ (12)
@@
 The negative distributions are denoted by $\mathcal{D}_{j}$, where $K$ represents the number of negative distributions.
-Note that the value of $K$ may vary depending on the training framework.
+Note that the $K$ may vary depending on the training framework.
 4.4 Analysis : Relation with Cosine Similarity Function
-Next, we establish the connection between the conventional cosine similarity function and our proposed DSF.
-We further explain why cosine similarity requires a temperature scaling term for effective optimization, whereas our DSF inherently eliminates the need for such adjustment.
+Next, we establish the connection between the conventional cosine similarity function and DSF.
 Theorem 1 (Relation between Divergence-based and Cosine Similarities).
-Let $Z_{i}=\{z^{1}_{i},\cdots,z^{m}_{i}\}$ and $Z_{j}=\{z^{1}_{j},\cdots,z^{m}_{j}\}$ are $m$ multi-view features.
-$\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ are the estimated von Mises-Fisher distribution from $Z_{i}$ and $Z_{j}$.
-Let $\text{sim}_{\text{cos}}(\cdot,\cdot)$ is cosine similarity and $\text{sim}_{\text{div}}(\cdot,\cdot)$ is the negative Kullback-Leibler divergence between two vMF distribution.
-If m=1, $\kappa_{1}$=$\kappa_{2}$ and $A_{p}(\kappa_{1})\cdot\kappa_{1}=1/\tau$ then
-$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\text{sim}_{\text{cos}}(z_{i},z_{j})/\tau+constant$ (12)
-Furthermore, InfoNCE Loss with $sim_{\text{div}}$ becomes equivalent to that using $sim_{\text{cos}}$ one:
-$L^{\text{div}}_{\text{InfoNCE}}=L^{\text{cos}}_{\text{InfoNCE}}$ (13)
+If $m=1$ and all concentration parameters $\kappa$ have the same value $\kappa^{*}$ and $A_{p}(\kappa^{*})\cdot\kappa^{*}=1/\tau$ then,
+$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\mathrm{sim}_{\mathrm{cos}}(z_{i},z_{j})/\tau+constant$ (13)
+Furthermore, InfoNCE Loss with $\mathrm{sim}_{\mathrm{div}}$ becomes equivalent to that using $\mathrm{sim}_{\mathrm{cos}}$ one:
+$L^{\mathrm{div}}_{\mathrm{InfoNCE}}=L^{\mathrm{cos}}_{\mathrm{InfoNCE}}$ (14)
 Proof.
-Proof is at Appendix A. ∎
-Theorem 1 says that when we have two views and kappa is set under certain conditions.
-Then our proposed divergence-based simiarity function becomes cosine simiarity function with some constant value.
-Furthermore, this leads to fact that InfoNCE loss becomes equivalent.
-Now, we explain the dependency of cosine similarity on temperature hyperparameter and independence of our DSF with respect to temperature hyperparameter.
+See Appendix A. ∎
+Theorem 1 states that when we have two views and $\kappa$ is set under certain conditions, DSF becomes cosine similarity with some constant value.
+Furthermore, this leads to the fact that InfoNCE losses become equivalent.
+We next provide an explanation of why cosine similarity requires a temperature scaling term for effective optimization, while DSF inherently eliminates the need for such adjustment.
 Proposition 1.
 Let $K$ be the number of negative pairs, and assume all negative similarities have the same value.
-Let $s^{+}$ and $s^{-}$ denote the positive and negative similarities, respectively.
+Let $s^{+}$ and $s^{-}$ denote the positive and negative similarity, respectively.
 Then, the InfoNCE loss becomes:
-$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(s^{+}-s^{-})}{\exp(s^{+}-s^{-})+K}\\ (s^{+}-s^{-})&\in\left\{\begin{matrix}[-\frac{2}{\tau},\frac{2}{\tau}]&\text{if }\text{sim}=\text{sim}_{\text{cos}}\\ (-\infty,\infty)&\text{if }\text{sim}=\text{sim}_{\text{div}}\end{matrix}\right.\end{split}$ (14)
+[…]
 In the optimal case for cosine similarity, we have $s^{+}=+1$ and $s^{-}=-1$, resulting in:
-$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(2/\tau)}{\exp(2/\tau)+K}\\ &=6.3196,\quad\tau=1.0,\,K=4096\\ &=4.3310,\quad\tau=0.5,\,K=4096\\ &=0.1706,\quad\tau=0.2,\,K=4096\\ &=0.0000,\quad\tau=0.1,\,K=4096\\ \end{split}$ (15)
-It has been theoretically shown (see Proposition1) that cosine similarity cannot represent a appropriate loss value without the use of a temperature term $\tau$.
-The core issue arises from the limited range of the similarity margin $(s^{+}-s^{-})$, which restricts the contrast between positive and negative pairs.
-To compensate for this limitation, the temperature parameter $\tau$ is introduced to amplify this margin, allowing for better gradient signals during training.
-The number of negative samples $K$ varies depending on the contrastive learning framework.
-For instance, in SimCLR and DCL, $K=2B-2=254$ when the batch size $B=128$, while in MoCo, $K$ is determined by the size of the memory queue, ranging from 4,096 to 65,536 depending on the dataset.
-As summarized in Table 2, the InfoNCE loss yields inappropriate values without the temperature parameter $\tau$ in all $K$ values.
-It says cosine similarity requires a carefully tuned temperature parameter across all frameworks to function effectively.
-In contrast, our proposed DSF expresses an appropriate loss without the need for such a temperature hyperparameter.
-See Figure 5 for $(s^{+}-s^{-})$ margin in our DSF experiment.
+[…]
+We show (see Proposition 1) that cosine similarity alone fails to produce appropriate loss values without a temperature term $\tau$, due to the limited similarity margin $(s^{+}-s^{-})$ between positive and negative pairs.
+The temperature parameter compensates for this by amplifying the margin.
+As shown in Table 2, InfoNCE loss with cosine similarity produces suboptimal values for all $K$ when $\tau$ is omitted, highlighting the need for careful tuning.
+In contrast, DSF achieves appropriate loss values without requiring such tuning.
+See Figure 5 for similarity in DSF.
 $L_{\text{InfoNCE}}$
@@
 Table 2: InfoNCE loss values under optimal similarity ($s^{+}=1$, $s^{-}=-1$) for varying temperature $\tau$ and the number of negative pairs $K$.
+The number of negatives $K$ depends on the framework: SimCLR uses K=2B2=254K=2B-2=254 when $B=128$, while MoCo sets $K$ by the memory queue size (4096–65536).
 Algorithm 1 Three Multi-View Methods
-Input: Instance $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
+Input: Data $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
 Parameter: Number of views $M=2m$.
@@
 1: Apply $M=2m$ augmentations to $x_{i}$, yielding $x_{i}^{1},\dots,x_{i}^{m}$ and $x_{B+i}^{1},\dots,x_{B+i}^{m}$
-2:  Extract features:
-$z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\quad z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\quad$
-for $\text{for }l=1,\dots,m$
-3:  if multi-view method = loss_avg:
-$L_{\text{loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
-4:  if multi-view method = fea_avg:
-$z_{i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$
-$L_{\text{fea\_avg}}=L_{\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
-5:  if multi-view method = DSF:  Estimate vMF distribution parameters as in Eq. (3)–(4):
-$(\mu_{i},\kappa_{i})\leftarrow\{z_{i}^{1},\dots,z_{i}^{m}\}$
-$(\mu_{B+i},\kappa_{B+i})\leftarrow\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}$
-6:   Compute divergence-based similarity:
-$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$
-$L_{\text{DSF}}=L_{\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
+2: Extract features: $z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\,z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\,\text{for }l=1,\dots,m$
+3:  if multi-view method = Loss_avg:   $L_{i,\text{Loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{i,\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
+4:  if multi-view method = Fea_avg:   $z_{i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$   $L_{i,\text{Fea\_avg}}=L_{i,\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
+5:  if multi-view method = DSF:   Estimate vMF distribution parameters as in  Eq.
+(3)–(4):    $(\mu_{i},\kappa_{i})\leftarrow g_{\text{est}}\left(\{z_{i}^{1},\dots,z_{i}^{m}\}\right)$    $(\mu_{B+i},\kappa_{B+i})\leftarrow g_{\text{est}}\left(\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}\right)$
+6:   Compute divergence-based similarity:    $\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$    $L_{i,\text{DSF}}=L_{i,\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
 5 Experiments
 5.1 Experimental Setup
-Every experiment is conducted with one or four NVIDIA GeForce RTX 3090.
-Dataset
-This study utilizes two benchmark datasets: CIFAR-10 and ImageNet-100.
-CIFAR-10 [Krizhevsky et al., 2009] consists of 60,000 color images of size $32\times 32$ pixels, divided into 50,000 training and 10,000 test images across 10 object classes.
-ImageNet-100 is a curated subset of the ImageNet dataset, comprising 100 randomly selected classes with approximately 1,300 images per class.
-It serves as a mid-scale benchmark for evaluating representation learning performance.
-These datasets are used to validate the proposed method on both small-scale and mid-scale visual recognition tasks.
-Model Arichitecture
-We primarily follow the MoCo v2 [Chen et al., 2020b] framework for contrastive learning.
-Throughout this paper, any reference to ”MoCo” refers to MoCo v2 unless otherwise specified.
-We adopt ResNet-18 as the encoder for experiments on CIFAR-10 The final fully connected layer projects features into a 128-dimensional embedding space.
-For large-scale evaluation, we extend our experiments to ImageNet-100 using MoCo v3 [Chen et al., 2021] with a Vision Transformer (ViT-S/16) backbone.
-This enables assessment of the scalability of the proposed method across both convolutional and transformer-based architectures.
-Data Augmentation
-For data augmentation, we apply a combination of RandAugment [Cubuk et al., 2020], ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in contrastive learning setups.
+Datasets.
+We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky, Hinton et al.
+2009) and ImageNet-100, a 100-class subset of ImageNet (Deng et al.
+2009).
+These datasets allow us to evaluate the proposed method on both small- and large-scale visual recognition tasks.
+We investigated transfer learning performance on the Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al.
+2009), FGVC Aircraft (Maji et al.
+2013), the Describable Textures Dataset (DTD) (Cimpoi et al.
+2014) and Oxford-IIIT Pets (Parkhi et al.
+2012).
+We adopt CIFAR-10-C (Hendrycks and Dietterich 2019) and ImageNet-100-C as corruption benchmarks to evaluate the robustness of the learned representations.
+The ImageNet-100-C dataset is constructed by selecting the 100 classes used in ImageNet-100 from the original ImageNet-C benchmark.
+For both datasets, we use only the corrupted samples with severity level 1 for evaluation.
+Model Architectures.
+We follow the MoCo v2 (Chen et al.
+2020b) framework, using ResNet-18 as the encoder for CIFAR-10, with a 128-dimensional projection head.
+For ImageNet-100, we adopt MoCo v3 (Chen, Xie, and He 2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
+Data Augmentation.
+For data augmentation, we apply a combination of RandAugment (Cubuk et al. 2020), ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in CL setups.
 5.2 Evaluations
-To evaluate the model’s performance, we conducted following tasks: k-Nearest Neighbors (kNN) and Linear Evaluation.
-Linear Evaluation assesses the representational quality of the encoder by attaching a classifier, while kNN measures similarity-based performance directly within the encoder’s embedding space.
+We evaluate the model using two standard protocols: k-Nearest Neighbors (kNN) and Linear Evaluation.
+The kNN evaluation measures similarity preservation in the embedding space, while linear evaluation assesses the representational quality via a trained classifier.
 kNN Evaluation.
-We evaluated the model with a kNN approach to measure the encoder’s capacity to retain similarity in the embedding space.
-Evaluation was conducted on CIFAR-10, ImageNet-100 using the output feature of the trained query encoder.
-The K value was set to 200, enabling an assessment of the encoder’s representations in a non-parametric, unsupervised manner.
-In our experiments, the proposed model achieved a kNN accuracy of 79.00% on the ImageNet-100 dataset, outperforming the baseline MoCo v3 model, which attained 74.08%.
-On the CIFAR-10 dataset, our model achieved 90.04% accuracy, compared to 88.44% with the MoCo baseline.
+Using $k=200$, we conduct kNN evaluation on CIFAR-10 and ImageNet-100 with features from the trained encoder.
+DSF method achieves 79.00% on ImageNet-100 and 90.04% on CIFAR-10, outperforming MoCo v3 baselines (74.08% and 88.44%, respectively) and other multi-view methods.
+See Table 1.
 Linear Evaluation.
-To further validate the effectiveness of our method, we evaluate the quality of the learned representations through linear evaluation.
-Specifically, after performing self-supervised pretraining on ImageNet-100 and CIFAR-10 using our proposed DSF framework, we freeze the encoder and train a linear classifier (a single-layer MLP) on top of the extracted features.
-On the ImageNet-100 dataset, our method achieved a linear evaluation accuracy of 85.66%, outperforming the baseline MoCo v3 model, which attained 81.52%.
-Similarly, on the CIFAR-10 dataset, our method achieved an accuracy of 91.21%, compared to 88.10% obtained by the MoCo baseline.
-These results demonstrate that DSF leads to stronger and more transferable representations across different datasets.
-(a) kNN Top-1 Accuracy with respect to number of view.
-(b) Linear Evaluation Top-1 Accuracy with respect to number of view.
+We freeze the pretrained encoder and train a linear classifier on top.
+DSF achieves 85.80% on ImageNet-100 and 91.21% on CIFAR-10, surpassing MoCo v3 results of 83.42% and 88.10% and also other multi-view methods.
+These results indicate that DSF produces more discriminative and transferable representations.
+See Table 1.
+Transfer Learning.
+Transfer learning is one of the ultimate goals of SSL.
+We evaluate the transfer performance on five natural image datasets under the fine-tuning setting, where all model weights are updated without freezing.
+More details are in Appendix D.3
+Robustness under Distribution Shift.
+To assess the robustness of the learned representations under distribution shift, we evaluate the models on common corruption benchmarks: CIFAR-10-C and ImageNet-100-C, following the protocol of (Hendrycks and Dietterich 2019).
+We use the mean Corruption Error (mCE) as the evaluation metric, measured using the model parameters obtained from linear evaluation on CIFAR-10 and ImageNet-100, respectively.
+No additional fine-tuning is performed during this evaluation.
+As shown in Table 1, DSF consistently outperforms both the baseline and other multi-view contrastive learning methods, achieving an mCE of 16.20% on CIFAR-10-C and 33.15% on ImageNet-100-C.
+These results indicate that the joint structure–capturing capability of DSF also enhances the model’s robustness to input corruptions.
+(a) kNN Top-1 Accuracy with respect to the number of view.
+(b) Linear Evaluation Top-1 Accuracy with respect to the number of view.
 Figure 4: Comparison of kNN and linear evaluation accuracies according to the number of views in three multi-view learning methods.
+Model Method Dataset (Top-1 Accuracy %)
+CIFAR10 CIFAR100 Pet Aircraft DTD Average
+MoCo v3 - 96.36 80.80 84.11 47.58 59.68 73.70
+MoCo v3 Loss_avg 96.76 81.55 83.11 47.17 58.67 73.45
+MoCo v3 Feature_avg 96.72 81.06 82.80 46.45 57.66 72.93
+MoCo v3 DSF 97.11 81.76 84.17 49.49 60.59 74.62
+Table 3: Comparison of transfer learning performance.
+Top-1 accuracy (%) on five downstream datasets using MoCo v3 representations pre-trained on ImageNet-100.
+Bold indicates the best performance, and underline indicates the second best.
 5.3 Same Resource Comparison
-Our method samples $M$ multiple augmented views from each image, which could potentially increase computational overhead.
-Therefore, it is essential to ensure that the proposed approach remains efficient under equivalent computational constraints.
-Computational Efficiency.
-Our method samples $M$ augmented views from each image instance, which could potentially increase computational overhead.
-To address this, we ensure fair comparisons by maintaining consistent computational budgets across all experiments.
-To ensure the practicality of our method, we carefully control and match computational budgets across all experiments.
-Specifically, we fix the product of batch size and number of views—i.e., $B\times M$—so that memory usage remains constant regardless of the number of views.
-For example, in our ImageNet-100 experiments (see Table 1), the baseline MoCo is trained with a configuration of $512\times 2$, whereas our method and other multi-view baselines use $128\times 8$, maintaining identical GPU memory usage.
-This configuration strategy is also consistently applied to CIFAR-10 experiments.
-Moreover, to equalize the overall training time, we proportionally reduce the number of training epochs for multi-view methods.
-As a result, the total runtime difference between our approach and baseline models is strictly limited to within 30 minutes—less than 4% and 10% of the entire training duration in the ImageNet-100 and CIFAR-10 setting.
-These measures ensure that the observed performance improvements are attributable to our proposed DSF method itself, rather than to increased computational resources.
+DSF samples multiple views per image, which may increase computational cost.
+To ensure fair comparison, we fix the product of batch size and number of views ($B\times M$) across all methods, maintaining equal GPU memory usage.
+For example, in ImageNet-100 experiments, MoCo v3 uses $512\times 2$, while DSF and other multi-view baselines use $128\times 8$, preserving memory usage.
+In CIFAR-10 experiments, MoCo uses $1024\times 2$, while DSF and other multi-view baselines adopt $256\times 8$.
+To align training time, we reduce epochs for multi-view methods proportionally.
+This keeps runtime differences within 1 hour for ImageNet-100 (under 7%) and 30 minutes for CIFAR-10 (under 10%), ensuring that performance gains arise from DSF itself rather than extra computation.
 5.4 More Resource Comparison
-5.4.1 More Time Cost: Faster Convergence with Divergence Similarity
-Self-supervised learning methods are known to require longer training times for convergence due to the nature of instance discrimination objectives and the inherent challenge of learning meaningful representations from unlabeled data.
-This issue becomes even more pronounced when dealing with complex datasets or limited batch sizes.
-To investigate convergence behavior, we extended the number of training epochs and compared model accuracy as a function of elapsed training time, as shown in Figure 3.
-While the baseline MoCo gradually improves over a long training schedule (55 hours at 8000 epochs), our proposed DSF reaches comparable accuracy within only 11 hours at 400 epochs, using merely 20% of the total training time.
-Specifically, compared to the Loss_avg and Feature_avg baselines trained for 800 epochs, our DSF reaches similar accuracy within just 400 epochs, achieving a 2× speedup in convergence.
-Moreover, we observe that the performance gap between our DSF and the two other multi-view baselines (Loss_avg, Feature_avg) remains consistently large, regardless of the total training time.
-This indicates that our method not only converges faster but also achieves a higher final accuracy that is not easily matched even with extended training.
-This result highlights that our divergence-based formulation facilitates more efficient optimization in self-supervised learning.
-Notably, all three multi-view methods considered in this comparison have identical per-epoch training time, ensuring a fair evaluation.
-5.4.2 More GPU Memory : Number of Views
-We sample multiple augmented views from each image; thus, it is important to investigate how the number of sampled views affects the quality of the learned representations.
-Prior studies [Tian et al., 2020, Shidani et al., 2024, Pototzky et al., 2022] have shown that increasing the number of augmentations generally leads to better representation quality.
-In our framework, the number of views directly impacts the accuracy of the estimated von Mises–Fisher (vMF) distribution.
-As the number of views increases, the estimated parameters $\mu$ and $\kappa$ become more reliable due to reduced estimation variance.
-Figure 4 presents the performance trends of three multi-view learning methods on both kNN and linear evaluation tasks, as the number of views $M$ increases.
-Across all view counts, DSF consistently outperforms the other two approaches.
-These results suggest that our divergence-based formulation provides stable and reliable improvements in multi-view learning settings.
+More Time Cost: Faster Convergence with Divergence Similarity
+Self-supervised learning often requires longer training to converge, due to the difficulty of learning meaningful representations without labels (Chen et al.
+2020a; Wang et al.
+2021a, b).
+To investigate convergence behavior, we compare model accuracies over elapsed training time (Figure 3).
+While MoCo requires 55 hours for 8000 epochs, DSF achieves similar accuracy in just 11 hours (400 epochs), using only 20% of the time.
+Compared to Loss_avg and Feature_avg, DSF also reaches similar accuracy with half the epochs and maintains a consistently higher performance throughout training.
+These results highlight the efficiency of DSF in optimizing self-supervised objectives.
+More GPU Memory: Consistent Superiority Across View Counts
+In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian, Krishnan, and Isola 2020; Shidani et al.
+2024; Pototzky, Sultan, and Schmidt-Thieme 2022).
+In DSF, more views lead to more accurate estimation of the vMF parameters $\mu$ and $\kappa$, due to reduced variance.
+As shown in Figure 4, DSF consistently outperforms other multi-view methods across different numbers of views, clearly demonstrating the superiority of DSF in multi-view representation learning.
 5.5 Temperature Free
-In numerous studies, the temperature parameter $\tau$ has been shown to be critical for achieving optimal performance in contrastive learning [Zhang et al., 2021, Zhang et al., 2022, Kukleva et al., 2023].
-In our method, however, $\tau$ is not required, as discussed in Proposition 1, and it can even degrade performance when introduced.
-To validate this, we conducted experiments over a wide range of temperature values, from 0.1 to 5.0.
-Interestingly, setting $\tau=1$ (i.e., using the raw similarity without scaling) consistently resulted in the best performance in Table 3.
-This suggests that our DSF is inherently numerically stable and does not require additional temperature tuning.
-Temperature $\tau$ 0.8 0.9 1 1.1
-kNN 91.2 91.3 91.8 91.4
-Linear Evaluation 92.0 92.3 92.5 92.0
-Table 3: Temperature Experiment. MoCo + DSF results in CIFAR10.
-5.6 Training Details
-To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch.
-We use AdamW as our optimizer.
-The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder.
-We choose batch size of 512, learning rate of 0.03.
-In our proposed method, a larger learning rate of 0.06 is used.
-GPU used for overall training is NVIDIA GeForce RTX 3090.
-6 Future Work
+While prior works (Zhang et al.
+2021, 2022; Kukleva et al.
+2023) emphasize the importance of the temperature parameter $\tau$ in CL, DSF does not require it (see Proposition 1) and even shows degraded performance when $\tau$ is introduced.
+As shown in Table 4, the best results are consistently achieved when $\tau=1$, indicating that DSF is inherently stable without temperature tuning.
+$\tau$ 0.1 0.9 1 1.1 5
+kNN 27.71 89.83 90.04 89.80 38.44
+Lin. Eval. 22.02 91.18 91.21 91.09 56.33
+Table 4: Temperature Experiment. Training MoCo with DSF under varying temperatures on CIFAR-10.
 Figure 5: Training dynamics of the similarity margin $(s^{+}-s^{-})$.
@@
 This plot shows the similarity margin of MoCo v3 with DSF ($128\times 8$) configuration on ImageNet-100.
-We use von Mises-Fisher distribution and Kullback-Leibler divergence for our proposed divergence-based similarity function.
-Although they are popular choice, there are plenty of available distributions and divergences.
-For example, in divergence, there are other divergences such as Rényi divergence, $\chi^{\alpha}$divergence, and Jensen–Shannon divergence—collectively known as the family of $f$-divergences could be explored.
-Investigating how these alternative divergences influence the representation quality and computational efficiency of contrastive methods remains a promising direction for future work.
+6 Conclusion
+In this work, we propose a novel similarity function, named the Divergence-based Similarity Function (DSF), which captures joint relationships among multiple features.
+Our method demonstrates strong performance under both equal and increased resource settings, surpassing baseline and other multi-view methods.
+In addition, it produces highly transferable and robust representations, as evidenced by its effectiveness in transfer learning and resilience to data corruptions.
+We also establish its connection to cosine similarity and demonstrate that DSF performs best with a fixed temperature of 1, eliminating the need for tuning.
+DSF is defined using the von Mises–Fisher distribution and Kullback–Leibler divergence.
+While effective, other $f$-divergences such as Rényi, $\chi^{\alpha}$, and Jensen–Shannon divergence remain unexplored, offering promising directions for future research.
+Beyond these avenues, a natural extension is a deeper theoretical study of DSF.
+In particular, examining its generalization behavior and convergence in high-dimensional regimes could provide stronger foundations for its empirical success.
+Such analysis may also clarify why DSF exhibits superior robustness and transferability compared to cosine similarity.
 Acknowledgements
-This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127) and MSIT/IITP(2021-0-00077)
+This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127, NRF-2016K2A9A2A13003815) and IITP(2021-0-00077).
+We would also like to thank Imseong Park, Hyunsung Bae, Jongyeon Lee and Jaesuk Kim, for their valuable feedback on an earlier version of this manuscript.
 Appendix A Appendix: Proof of Theorem 1
@@
 Thus, Feature_avg does not model joint structure beyond pairwise relationships.
-Appendix C Stabilization of the Concentration Parameter $\kappa$
-To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Equation 5.
+Appendix C Further Explanation of DSF
+C.1 Definitions
+In here, we introduce the definition of $A_{p}(\cdot)$ and $I_{p}(\cdot)$ in Eq.
+(3) and Eq.
+(LABEL:eq:kld) respectively.
+Here, $\Gamma$ denotes the gamma function.
+$I_{p}(\cdot)$ is Modified Bessel function of the first kind with order $p$.
+$A_{p}(\cdot)$ is a fraction of two Modified Bessel function of the first kind.
+$\begin{split}&I_{p}(\kappa)=\sum_{m=0}^{\infty}\frac{1}{m!\Gamma(m+p+1)}(\frac{\kappa}{2})^{2m+p},\\ &A_{p}(\kappa)=\frac{I_{\frac{p}{2}}(\kappa)}{I_{\frac{p}{2}-1}(\kappa)}\end{split}$ (20)
+C.2 Second Term Calculation
+In Eq.
+(LABEL:eq:kld), the second term cannot be evaluated in closed form using the modified Bessel function $I_{p}(\cdot)$.
+Therefore, we adopt an asymptotic approximation as given in (DLMF , Eq.
+(10.41.3)).
+[…]
+Using this approximation, we compute the second term as follows, truncating the series at $k=5$.
+$\begin{split}&\log\frac{I_{\frac{p}{2}-1}(\kappa_{1})}{I_{\frac{p}{2}-1}(\kappa_{0})}=\log\frac{I_{v}(vz_{1})}{I_{v}(vz_{0})}\\ &=v(\eta_{1}-\eta_{0})+\frac{1}{4}(\frac{1+z_{0}^{2}}{1+z_{1}^{2}})+\log\sum_{k=0}^{\infty}\frac{U_{k}(x_{1})}{v^{k}}\cdot\frac{1}{{\sum_{k=0}^{\infty}\frac{U_{k}(x_{0})}{v^{k}}}}\\ &v=\frac{p}{2}-1\end{split}$ (22)
+C.3 Stabilization of the Concentration Parameter $\kappa$
+To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Eq.
+(5).
 However, even this approximation can lead to excessively large values of $\kappa$ as training progresses.
 To mitigate this, we propose two stabilization techniques.
-Figure 6 illustrates the original approximated $\kappa$ values computed from Equation 5, alongside the stabilized values obtained after applying our proposed techniques.
+Figure 6 illustrates the original approximated $\kappa$ values computed from Eq. (5), alongside the stabilized values obtained after applying our proposed techniques.
 Figure 6: Consistency by Kappa Normalization.
-$\kappa$ value is calculated as approximated formulation as (5).
+$\kappa$ value is calculated as approximated formulation as Eq.
+(5).
 $x$-axis is $\bar{R}$ and $y$-axis is calculated $\kappa$.
@@
 Without these two techniques, $\kappa$ value is going excessively large.
+Appendix D Experimental Details
+D.1 Training Details
+Every experiment is conducted with one or four NVIDIA GeForce RTX 3090.
+To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch.
+We use AdamW as our optimizer.
+The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder.
+We choose batch size of 512, learning rate of 0.03.
+In DSF, a larger learning rate of 0.06 is used.
+GPU used for overall training is NVIDIA GeForce RTX 3090.
+D.2 Transfer Learning
+We conduct transfer learning experiments following the protocol in (Chen et al.
+2020a).
+We report top-1 accuracy for CIFAR-10, CIFAR-100 and the Describable Textures Dataset (DTD); and mean per-class accuracy for FGVC Aircraft and Oxford-IIIT Pets.
+D.3 Robustness under Distribution Shift
+We adopt the corruption methodology introduced by  (Hendrycks and Dietterich 2019).
+The ImageNet-100-C dataset is constructed by extracting the 100 selected classes from the original ImageNet-C benchmark.
+We evaluate the models using the mean Corruption Error (mCE) metric on corrupted images with severity level 1.
+The evaluation covers all 18 corruption types provided in the benchmark: gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, brightness, contrast, elastic transform, pixelate, jpeg compression, speckle noise, gaussian blur, spatter, and saturate.
+(a) Three terms and KL divergence value of positive pairs.
+(b) Three terms and KL divergence value of negative pairs.
+Figure 7: (Top) Positive pairs.
+(Bottom) Negative pairs.
+Positive pairs maintain KL divergence values close to 0, reflecting their similarity, while negative pairs exhibit larger divergence values, reflecting increasing dissimilarity.

### v4

@@
 Furthermore, we establish a connection between DSF and cosine similarity, and demonstrate that, unlike cosine similarity, DSF operates effectively without the need for tuning a temperature hyperparameter.
+Code and pretrained models are available at https://github.com/Jeon789/DSF.
 1 Introduction
@@
   • Loss-level approaches: Averaging pairwise losses computed between all available combinations of views (Shidani et al. 2024).
-  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky, Sultan, and Schmidt-Thieme 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
+  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky et al. 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
 Figure 1: Illustration of divergence-based similarity function: given multiple augmented views of the image, their extracted feature vectors (colored dots) lie on the hypersphere and are used to construct probability distributions (colored regions) over the unit hypersphere.
@@
 We estimate a probability distribution on the unit hypersphere and compute similarity via negative divergence between distributions, see Figure 1.
-We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher, Lewis, and Embleton 1993; Kitagawa and Rowley 2022; Scott, Gallagher, and Mozer 2021; Du et al.
+We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher et al.
+1993; Kitagawa and Rowley 2022; Scott et al.
+2021; Du et al.
 2024).
 vMF distribution is parameterized by mean direction $\mu$ and concentration $\kappa$, which respectively represent the central tendency and joint structure of features.
-Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord, Li, and Vinyals 2018), linking pairwise and multi-view contrastive frameworks.
+Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord et al.
+2018), linking pairwise and multi-view contrastive frameworks.
 Our main contributions are:
@@
 SimCLR (Chen et al.
-2020a) leverages the InfoNCE loss (Oord, Li, and Vinyals 2018) to pull together representations of two augmented views while pushing apart all other samples.
+2020a) leverages the InfoNCE loss (Oord et al.
+2018) to pull together representations of two augmented views while pushing apart all other samples.
 MoCo (He et al.
@@
 2022) decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable and high performance training with smaller batches and without a momentum encoder.
-MoCo v3 (Chen, Xie, and He 2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al.
+MoCo v3 (Chen et al.
+2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al.
 2020).
@@
 The second category is a method that deals with view representations at the feature level before computing the final objective.
-FastSiam (Pototzky, Sultan, and Schmidt-Thieme 2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al.
+FastSiam (Pototzky et al.
+2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al.
 2024) adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost.
@@
 Datasets.
-We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky, Hinton et al.
+We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky et al.
 2009) and ImageNet-100, a 100-class subset of ImageNet (Deng et al.
@@
 These datasets allow us to evaluate the proposed method on both small- and large-scale visual recognition tasks.
-We investigated transfer learning performance on the Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al.
+We investigated transfer learning performance on the Food-101 dataset (Bossard et al.
+2014), CIFAR-10 and CIFAR-100 (Krizhevsky et al.
 2009), FGVC Aircraft (Maji et al.
@@
 2020b) framework, using ResNet-18 as the encoder for CIFAR-10, with a 128-dimensional projection head.
-For ImageNet-100, we adopt MoCo v3 (Chen, Xie, and He 2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
+For ImageNet-100, we adopt MoCo v3 (Chen et al.
+2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
 Data Augmentation.
@@
 More GPU Memory: Consistent Superiority Across View Counts
-In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian, Krishnan, and Isola 2020; Shidani et al.
-2024; Pototzky, Sultan, and Schmidt-Thieme 2022).
+In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian et al.
+2020; Shidani et al.
+2024; Pototzky et al.
+2022).
 In DSF, more views lead to more accurate estimation of the vMF parameters $\mu$ and $\kappa$, due to reduced variance.
@@
 (LABEL:eq:kld), the second term cannot be evaluated in closed form using the modified Bessel function $I_{p}(\cdot)$.
-Therefore, we adopt an asymptotic approximation as given in (DLMF , Eq.
+Therefore, we adopt an asymptotic approximation as given in (DLMF 2024, Eq.
 (10.41.3)).